OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)?

OpenStudy (anonymous):

find slope=(y2-y1)/(x2-x1) then put the slope value in this equation y-y1=m(x-x1) where m=slope